OpenStudy (anonymous):

Solve the equation for y. 5x – 7y = 10

OpenStudy (anonymous):

ok ! we can say : -7y=-5x+10 right ? @zobobozo

OpenStudy (anonymous):

truu

OpenStudy (anonymous):

then how you can take out -7 from y ?

OpenStudy (anonymous):

O_o I got x-10=7/5y

OpenStudy (anonymous):

-7y=-5x-7

OpenStudy (anonymous):

and you know we can divided all of equation in a number .

OpenStudy (anonymous):

yep

OpenStudy (ranga):

5x – 7y = 10 Solve the equation for y means we have to isolate y. Add -5x to both sides: -7y = 10 - 5x divide both sides by -7 y = -10/7 + 5/7x y = 5/7x - 10/7